Identify and trace the conicoid y2 +z2 = x. Describe its sections by the planes x = 0, y = 0 and z = 0


1
Expert's answer
2021-06-14T17:56:57-0400

"y^2+z^2=x"


Now we can see that our surface is spherical ellipsoid.


Denominators are equal. It means that our ellipsoid is not a ellipsoid of revolution.


Let "y=0", then "z^2 = x"

And let "z=0", then "y^2=x"

"z^2 = x" and "y^2=x" principal parabolas.



Let's look on sections.


"\implies" "y=0"

"\implies" "z^2 = x" ...................Parabola with its axis along x - axis.



"\implies"  "z=0"

"\implies""y^2 = x" .................Parabola with its axis along x - axis.




"\implies" "x=0" 

"\implies""y^2 + z^2 = 0" 


"y^2+z^2=x"


Now we can see that our surface is spherical ellipsoid.


Denominators are equal. It means that our ellipsoid is not a ellipsoid of revolution.


Let "y=0", then "z^2 = x"

And let "z=0", then "y^2=x"

"z^2 = x" and "y^2=x" principal parabolas.



Let's look on sections.


"\implies" "y=0"

"\implies" "z^2 = x" ...................Parabola with its axis along x - axis.



"\implies"  "z=0"

"\implies""y^2 = x" .................Parabola with its axis along x - axis.




"\implies" "x=0" 

"\implies""y^2 + z^2 = 0" ...........It represents x- axis.


Thecurve is shown below: